# Across Text and Vision [Version 0.1]

### v1

@@ Abstract
Chain-of-Thought (CoT) reasoning has been widely adopted to enhance Large Language Models (LLMs) by decomposing complex tasks into simpler, sequential subtasks. However, extending CoT to vision-language reasoning tasks remains challenging, as it often requires interpreting transitions of visual states to support reasoning. Existing methods often struggle with this due to limited capacity of modeling visual state transitions or incoherent visual trajectories caused by fragmented architectures.
To overcome these limitations, we propose Uni-CoT, a Unified Chain-of-Thought framework that enables coherent and grounded multimodal reasoning within a single unified model. The key idea is to leverage a model capable of both image understanding and generation to reason over visual content and model evolving visual states. However, empowering a unified model to achieve that is non-trivial, given the high computational cost and the burden of interleaved image-text training. To address this, Uni-CoT introduces a novel two-level reasoning paradigm: A Macro-Level CoT for high-level task planning, and A Micro-Level CoT for subtask execution. This hierarchical design significantly reduces the computational overhead. Furthermore, we introduce a structured training paradigm that combines interleaved image-text supervision for macro-level planning with multi-task auxiliary objectives for micro-level execution. Together, these innovations allow Uni-CoT to perform scalable, stable, and high-fidelity multi-modal reasoning over dynamic visual states. Furthermore, thanks to our design, all experiments can be efficiently completed using only 8 A100 GPUs with 80GB VRAM each. Experimental results on reasoning-driven image generation benchmark (WISE) and editing benchmarks (RISE and KRIS) indicates that Uni-CoT demonstrates state-of-the-art performance and strong generalization, establishing Uni-CoT as a promising solution for multi-modal reasoning. Project Page and Code: https://sais-fuxi.github.io/projects/uni-cot/
Figure 1: Interpretable multi-modal reasoning trajectory of Uni-CoT. Uni-CoT extends Chain-of-Thought (CoT) reasoning to the multi-modal domain, enabling a unified model to perform coherent, grounded, and step-by-step reasoning across text and images for complex multi-modal tasks.
@@ 1 Introduction
Chain-of-Thought (CoT) reasoning [38] has emerged as a powerful paradigm for enhancing the performance of Large Language Models (LLMs) on complex tasks [10, 5, 23], by explicitly generating intermediate reasoning steps to enable step-by-step problem solving. Motivated by its success, recent works [50, 24, 53] have spent efforts to extend CoT to multi-modal settings, aiming to equip Multi-modal Large Language Models (MLLMs) with the capacity to handle vision-language reasoning tasks such as visual question answering [43], image editing with textual instructions [11], and embodied planning in interactive environments [25].
Prior works [49, 46, 9, 17, 41] explored utilizing reinforcement learning (RL) to enhance the text-based reasoning capabilities of MLLMs. While these approaches have shown promise, they often struggle in complex scenarios that require dynamic visual reasoning, such as geometric problems requiring diagram updates, visual puzzles demanding iterative image edits, or performing object rearrangement tasks involving spatial and causal understanding [39, 4]. A key reason of their failures possibly stems from the fundamental differences between textual and multi-modal reasoning: unlike purely text-based tasks, human multi-modal reasoning involves interpreting dynamic visual state transitions [48, 1], including object motion, spatial layouts, and visual causality, that naturally assist human reasoning but are inherently difficult to model through discrete symbolic representations alone.
@@
Motivated by this, recent studies [14, 33, 15, 16, 31, 12] have explored programmatic visual manipulations, such as cropping, line plotting, and depth or segmentation estimation, to simulate visual transitions. While these methods capture certain dynamics within images, they fall short in modeling the structural changes critical for complex tasks like sequential image editing or goal-directed navigation, where visual states evolve significantly over time. To address these limitations, several approaches [44, 52, 13] adopt modular frameworks that combine MLLMs with image/video generators to achieve dynamic visual state transitions. However, the disjointed gradient flow between the reasoning and generation components in these works leads to fragmented reasoning trajectories and visually inconsistent or implausible visual state transitions, ultimately hindering the system’s ability to perform coherent and grounded multi-modal reasoning.
@@
To address these limitations, we propose Uni-CoT, a Unified Chain-of-Thought framework, designed to support structural visual transitions for coherent multi-modal reasoning. Our core idea is to leverage a unified model [7], capable of both image understanding and generation, to support reasoning grounded in visual content and to model dynamic visual state transitions. This design is motivated by the intuition that using an unified model [34, 42, 3, 35] for both reasoning and generation naturally minimize the discrepancy between reasoning trajectories and visual transitions, thereby ensuring coherence in multi-modal inference. In addition, a unified architecture enables seamless implementation of end-to-end fine-tuning and reinforcement learning, facilitating higher coherence and performance ceilings for multi-modal reasoning tasks.
However, enabling the unified model to achieve multi-modal reasoning is not intuitive due to two significant challenges: (1) Computational Complexity: Unlike text-only reasoning within around 300 tokens per step, multi-modal reasoning jointly generate text and visual intermediate results, requiring around 10,000 tokens per step (see Sec. 2). This dramatic increase in sequence length significantly raises the computation and storage overhead of self-attention, limiting scalability and and even hindering the convergence of training. (2) Training Complexity: Interleaving image and text generation introduces significant optimization challenges due to mismatched learning dynamics and loss scales. Balancing cross-entropy for text with MSE for images often leads to instability, and reinforcement learning further amplifies this by requiring modality-aware rewards and careful optimization to maintain training stability.
To address computational complexity, we propose a novel two-level multi-modal reasoning framework comprising a Macro-Level CoT for high-level planning and a Micro-Level CoT for subtask execution. Specifically, the Macro-Level CoT performs global task planning by decomposing complex problems into simpler subtasks and aggregating their outcomes, while deliberately abstracting away execution detail of subtask. In contrast, the Micro-Level CoT focuses on solving individual subtasks while filtering out irrelevant information. To further simplify the reasoning process, we model micro-level reasoning as a Markov Decision Process (MDP), where each step depends only on the preceding state and current subtask instruction. With this hierarchical design, we decompose long reasoning trajectories to multiple short segments, thereby minimizing unnecessary token interactions and significantly reducing the computation complexity of self-attention layers.
@@
Building upon this hierarchical design, the training of multi-modal reasoning can be naturally decomposed into two streamlined components: (1) Interleaved text-image supervision for the Macro-Level CoT branch, which guides the model in learning global planning strategies and final result synthesis from high-level trajectories; and (2) Multi-task auxiliary learning for the Micro-Level CoT branch, which supervises the MDP-style execution via a set of simplified objectives such as action generation, next-state prediction, and reward estimation. This decoupled training paradigm enables effective supervision at both global and local levels, facilitating efficient and scalable learning for complex multi-modal reasoning tasks.
@@ 2 Preliminary: The Unifed Model - BAGEL
To enable unified chain-of-thought (CoT) reasoning across both image and text modalities, our framework builds upon BAGEL (Scalable Generative Cognitive Model), a recently released open-source foundation model that supports joint vision-language understanding and generation. We briefly introduce the core elements of BAGEL in following sections.
@@
BAGEL adopts a unified, decoder-only transformer architecture featuring a Mixture-of-Transformer-Experts [21] design. Two experts are instantiated: one dedicated to understanding and the other to generation. Both experts operate over shared multi-modal token sequences through a unified self-attention mechanism, allowing flexible and lossless fusion across modalities without introducing task-specific bottlenecks.
Specifically, BAGEL integrates two modality-specific visual encoders:
  • Vision Transformer (ViT) encoder for semantic-level understanding, which transforms an input image into approximately 4,900 tokens. It is initialized from a SigLIP2 [36] model and supports arbitrary aspect ratios via NaViT [6]-style positional encoding.
  • Variational Autoencoder (VAE) from FLUX [20] for pixel-level generation, which encodes an image into a latent grid of $\times64$ with 16 channels, resulting in 4,096 latent tokens.
The expert routing mechanism adopts hard gating: the understanding expert is activated for text and ViT tokens, while the generation expert exclusively processes VAE tokens. This dual-pathway design supports both fine-grained visual synthesis and high-level semantic grounding within a single, autoregressive modeling framework.
@@
  • Cross-Entropy Loss for autoregressive token prediction over text tokens:
@@
    where $x_{i}$ denotes the target token, $\hat{x}_{i}$ is the predicted token and $C$ is the number of token classes, which may be textual or visual.
  • Mean Squared Error Loss. For denoising-based generation under the Rectified Flow [22] paradigm, given a clean latent $\mathbf{x}_{0}$ and a Gaussian noise sample $\mathbf{x}_{1}$, we construct the noisy latent $\mathbf{x}_{t}$ via linear interpolation:
@@
    The model is then trained to predict the velocity field using the Mean Squared Error (MSE) loss:
@@
    Then the total training loss is a weighted combination of both objectives:
@@
    with empirically tuned weights $\lambda_{\text{CE}}=0.25$ to ensure training stability.
@@
At inference time, BAGEL operates autoregressively over interleaved multi-modal token sequences:
  • Understanding tasks (e.g., visual question answering, multi-modal reasoning): BAGEL consumes ViT-encoded image tokens and text tokens, and outputs the next text token via standard next-token prediction.
  • Generation tasks (e.g., text-to-image synthesis, image editing): BAGEL predicts VAE latent tokens conditioned on prompts or reference images and denoises them via Rectified Flow.
This unified inference mechanism enables BAGEL to handle diverse multi-modal tasks within a single autoregressive interface.
@@
As discussed above, while BAGEL provides a unified modeling backbone, it faces significant computational bottlenecks when applied to step-wise multi-modal reasoning. Unlike text-only Chain-of-Thought (CoT), where each reasoning step typically involves 512–1,024 tokens, multi-modal CoT often requires both image understanding and image generation within each step, introducing substantial overhead: Generating an image via VAE incurs approximately 4,096 tokens and Encoding an image via ViT for understanding introduces an additional 4,900 tokens. This results in nearly 9,000 visual tokens, in addition to around 1,000 text tokens, per reasoning step. As the number of steps increases, the overall training and inference costs become prohibitively expensive. Moreover, this token-intensive formulation imposes significant challenges on model optimization, often hindering convergence and limiting generalization, particularly in complex tasks that require long and compositional reasoning chains.
3 Methodology
Figure 2: Overview of the Uni-CoT framework. Uni-CoT operates via two complementary branches: Macro-Level CoT decomposes a complex task into simpler subtasks and synthesizes their outcomes to derive the final answer, while keeping intra-subtask reasoning implicit to reduce learning and computational overhead; Micro-Level CoT formulates each subtask as a Markov Decision Process (MDP), where each reasoning step (a text-image pair) is a discrete node, and transitions depend solely on the previous step and current subtask prompt.
Figure 3: Uni-CoT masking strategy for hierarchical reasoning.(a) Macro masked attention scheme: Only the system prompt, high-level planning outputs, and final subtask outcomes are visible to the model. Intermediate reasoning traces—such as textual rationales or image edits produced during subtask execution—are fully masked. (b) Micro masked attention scheme: Each self-check (self-reflection) step can only attend to the immediate previous state (e.g., the last image-text pair) and the current subtask instruction.
To address the inherent complexity of multi-modal reasoning, Uni-CoT adopts a hierarchical reasoning framework inspired by human cognitive strategies. Specifically, the reasoning process is divided into two levels: a Macro-Level CoT, which performs high-level task planning and summarization by decomposing complex problems into a sequence of simpler subtasks and aggregating their outcomes to synthesize the final result; and a Micro-Level CoT, which focuses on subtask execution through iterative, feedback-driven reasoning to ensure stable and reliable intermediate outputs. This two-tiered architecture promotes efficient reasoning, improved generalization, and enhanced interpretability across diverse multi-modal tasks.
We detail the planning strategies of Macro-Level CoT in Sec. 3.1 and the subtask execution mechanism of Micro-Level CoT in Sec. 3.2, followed by an explanation of the training procedures in Sec. 3.3.
@@ 3.1 Macro-Level CoT: Planning Strategies
Given a complex task, humans typically begin by outlining one or more abstract pathways toward the goal, rather than considering every low-level detail from the outset. Inspired by this cognitive behavior, a central function of the Macro-Level CoT is to formulate a global plan that decomposes the original task into a set of simpler, tractable subtasks.
@@
Specifically, drawing inspiration from [8, 45, 28, 47], we propose three cognitively motivated planning mechanism:
  1. 1.
    Sequential Decomposition Mechanism: Humans often tackle complex tasks by addressing intermediate goals in a step-by-step manner. Analogously, we define a Sequential Decomposition strategy, wherein a task is split into a fixed sequence of subtasks. This structured progression simplifies the overall reasoning path and improves traceability.
  2. 2.
    Parallel Decomposition Mechanism: In collaborative scenarios, tasks are frequently divided into independent components that can be solved concurrently. Motivated by this, we propose a Parallel Decomposition strategy, which enables the model to reason over multiple subtasks in parallel, leveraging task modularity to improve efficiency.
  3. 3.
    Implicit Planning via Progressive Refinement Mechanism: In uncertain or dynamic environments, such as navigating a maze, human often forego rigid plans and instead refine their decisions iteratively. To emulate this behavior, we propose a Progressive Refinement strategy, wherein the model incrementally refines its plan and revise earlier steps if inconsistencies arise, thereby supporting adaptive and flexible reasoning.
These strategies empower the Macro-Level CoT to effectively address the “what-to-do” aspect of reasoning.
In addition to high-level task planning, the Macro-Level CoT is also responsible for summarizing and analyzing the outputs of individual subtasks and synthesizing them into a coherent final answer, as illustrated in Figure 2. Crucially, we treat the internal reasoning process within each subtask as implicit from the macro perspective, these fine-grained execution detail are abstracted away and handled entirely by the Micro-Level CoT. This deliberate abstraction allows the macro planner to operate from a global vantage point, focusing solely on the structural decomposition and overall reasoning trajectory without being entangled in the intricacies of subtask execution.
@@
To enforce this abstraction during both training and inference, we introduce a macro masked-attention scheme, depicted in Figure 3 (a). This masking mechanism selectively reveals only the system prompt, macro-level planning outputs, and the final subtask outcomes. All intermediate reasoning traces, such as textual rationales and image modifications produced during micro-level subtask execution, are completely masked out. By restricting the visible context to only high-level information, this scheme encourages the model to reason at the task-structural level and prevents it from overfitting to low-level execution details. As a result, the Macro-Level CoT learns to maintain a top-down reasoning perspective, fostering modularity and scalability in multi-modal Chain-of-Thought reasoning.

### v2

# Text and Vision [Version 0.2: For Nano-Banana Like Generation]

@@ Abstract
Chain-of-Thought (CoT) reasoning has been widely adopted to enhance the performance of Large Language Models (LLMs) on complex tasks by step-by-step problem solving. However, extending CoT to complex vision-language tasks remains challenging, as it often requires interpreting visual change to assist reasoning. Existing methods often struggle with this due to limited capacity of modeling visual state transitions or incoherent visual trajectories caused by fragmented architectures.
To overcome these limitations, we propose Uni-CoT, a Unified Chain-of-Thought framework that enables coherent visual state transition and grounded text reasoning for multi-modal tasks. Our key idea is to leverage a unified model, capable of both image understanding and generation, to reason over visual content and model visual transitions. However, empowering a unified model to achieve that is non-trivial due to the high computational cost and instable training. To address this, Uni-CoT introduces a novel two-level reasoning paradigm: A Macro-Level CoT for high-level task planning, and A Micro-Level CoT for subtask execution. This hierarchical design significantly reduces computational overhead by localizing reasoning dependencies. Furthermore, we introduce a structured training paradigm that combines multi-modal supervision for macro-level CoT with multi-task supervision for micro-level CoT. Together, these innovations allow Uni-CoT to perform scalable and stable multi-modal reasoning with coherent visual transitions. Furthermore, with our designs, our model can be efficiently trained only by 8 A100 GPUs. Experimental results on reasoning-driven image generation benchmark (WISE) and editing benchmarks (RISE and KRIS) indicates that Uni-CoT demonstrates state-of-the-art performance and strong generalization, establishing Uni-CoT as a promising solution for multi-modal reasoning. Project Page and Code: https://sais-fuxi.github.io/projects/uni-cot/
Figure 1: The multi-modal reasoning trajectory of Uni-CoT. Uni-CoT extends Chain-of-Thought to the multi-modal domain, enabling a unified model to perform coherent, grounded, and step-by-step reasoning across text and images. More results refer to Figure S3.
@@ 1 Introduction
Chain-of-Thought (CoT) [68] is an approach that enhances the performance of Large Language Models (LLMs) on complex tasks by imitating the way humans solve problems. It works by encouraging LLMs to produce explicit intermediate reasoning steps, allowing them to tackle complex tasks in a step-by-step manner. Motivated by its success [20, 14, 41], recent works [84, 42, 87] have paid efforts to extend CoT to multi-modal domain, aiming to equip Multi-modal Large Language Models (MLLMs) with the reasoning capacity to handle complex vision-language tasks, such as complex visual question answering [75], image editing [21], and embodied planning [43].
Prior efforts [82, 78, 19, 30, 72, 65, 57] have employed reinforcement learning (RL) to enhance the text-based reasoning capabilities of multi-modal large language models (MLLMs). While effective in text-centric applications, these approaches reveal notable limitations in domains that require strong visual reasoning, such as geometric analysis or visual navigation [70, 12]. In fact, many of these visual reasoning tasks can be easily solved by middle-school students, underscoring a fundamental gap between MLLMs and human cognitive abilities. Existing studies [17, 60] suggest that this human advantage stems from the natural integration of visual state transitions into reasoning. Such skills are cultivated early in multi-modal environments, where even simple activities like cooking require interpreting textual instructions while monitoring visual cues. However, this capacity remains challenging for current MLLMs to replicate, highlighting the need for future models to explicitly embed visual state transitions into their reasoning processes to unlock more general multi-modal reasoning abilities.
@@
Motivated by this gap, recent studies [27, 58, 28, 29, 55, 22] have explored programmatic visual manipulations (e.g., cropping, line plotting) to approximate visual state transitions. While effective at capturing local image dynamics, these methods fall short in modeling the structural visual changes needed for complex tasks such as goal-directed navigation or object rearrangement. To overcome this, other approaches [76, 86, 25] couple MLLMs with image or video generators, thereby enabling larger-scale transitions. In parallel, recent works [26, 32] leverage external reward models to guide generators through iterative visual refinements. However, the loose integration between reasoning and generation in these frameworks often results in fragmented reasoning flows and inconsistent transitions, ultimately limiting coherent multi-modal reasoning.
@@
To address these limitations, we propose Uni-CoT, a Unified Chain-of-Thought framework, designed to support structural visual transitions and coherent text understanding for multi-modal tasks. Our preliminary idea is to leverage a unified model [15, 59, 74, 11, 61], capable of both image understanding and generation, to support reasoning grounded in visual content and modeling dynamic visual state transitions. This design is motivated by the intuition that using a single model for both reasoning and generation naturally minimize the discrepancy between reasoning trajectories and visual transitions, thereby ensuring coherence in multi-modal reasoning. In addition, a unified architecture enables seamless implementation of end-to-end fine-tuning and reinforcement learning, facilitating higher coherence and performance ceilings for multi-modal tasks.
However, enabling a unified model to perform multi-modal reasoning is non-trivial due to two major challenges: (1) Computational Burden: Unlike text-only reasoning which consumes roughly= 300 tokens per step, multi-modal reasoning need to jointly produce textual and visual intermediates, requiring up to 10,000 tokens per step (see Sec. 2). This dramatic increase in sequence length greatly amplifies both computational and storage overhead. (2) Training Instability: Longer sequences demand the ability to model long-range dependencies, making optimization more difficult. In addition, interleaved image-text generation exacerbates instability, as the two modalities exhibit mismatched learning dynamics. Reinforcement learning further compounds these issues, requiring carefully designed modality-aware rewards and optimization strategies.
Fortunately, humans face similar difficulties when solving multi-modal complex tasks and offer insights into potential solutions. In real-world, complex multi-modal reasoning places heavy demands on working memory and often becomes challenging when integrating different types of information with multiple input channels [47]. For instance, solving a large jigsaw puzzle through brute-force trial quickly becomes intractable as the search space grows exponentially, overwhelming cognitive capacity. A well-established strategy for overcoming such challenges is the hierarchical organization of reasoning [5, 45]. At a higher level, abstract reasoning generates a global plan that decomposes the problem into manageable subtasks; at a lower level, situated reasoning focuses on executing these subtasks with minimal interference. Returning to the puzzle example, experienced solvers often segment the board into smaller regions based on color or object similarity as a high-level planning and then assemble small areas one by one as low-level inference. The successful assembly of local areas provides intermediate feedback, progressively guiding the completion of the entire puzzle. This dual-level structure shortens reasoning trajectories and reduces cognitive load, enabling humans to sustain coherent reasoning over long horizons without being overwhelmed by complexity.
Inspired by this human cognitive ability, we propose a macro–micro hierarchical reasoning framework to mitigate the complexity of multi-modal reasoning. At the macro-level CoT, our model performs global planning by decomposing a complex task into manageable subtasks and synthesize the final result only according to subtasks’ results. At the micro-level CoT, the model focuses on solving each subtask with minimal interference from irrelevant context. To minimize the computational burden, the micro-level CoT is further formulated as a Markov Decision Process (MDP). By transforming a lengthy and cognitively demanding reasoning trajectory into a series of structured reasoning blocks, our hierarchical design reduces redundant token interactions in CoT, thereby substantially lowering computational complexity and improving training stability.
@@
Building on this hierarchical design, we further decompose multi-modal CoT learning into two complementary components: (1) Macro-Level CoT Modeling. The model is directly refined on interleaved text–image content to acquire the ability of global planning and final result synthesis. (2) Micro-Level CoT Modeling. The micro-level reasoning is further decomposed into an MDP-style process, where we design four auxiliary tasks, such as action generation and reward estimation, to facilitate effective and efficient learning. This decoupled paradigm provides supervision at both global and local levels, thereby enabling scalable and efficient training for complex multi-modal reasoning tasks. Moreover, to enhance optimization stability, we introduce a node-based reinforcement learning strategy. With these training designs, Uni-CoT can be trained efficiently on only 8 A100 GPUs. Extensive experiments on reasoning-driven image generation and editing benchmarks show that Uni-CoT delivers state-of-the-art performance, establishing it as a promising solution for multi-modal reasoning.
@@ 2 Preliminary: The Unifed Model - BAGEL
To enable unified chain-of-thought (CoT) reasoning across both image and text modalities, our framework builds upon BAGEL (Scalable Generative Cognitive Model), a recently released open-source unified model that supports joint vision-language understanding and generation. We briefly introduce the core elements of BAGEL in below.
@@
BAGEL adopts a unified decoder-only Transformer architecture and incorporates a Mixture-of-Transformer-Experts [36] design. It is initialized from Qwen-2.5 [2]. Bagel consists of two experts: one dedicated to understanding and the other to generation. Both experts operate over shared multi-modal token sequences through a unified self-attention mechanism, allowing flexible and lossless fusion across modalities without introducing task-specific bottlenecks. Specifically, BAGEL integrates two modality-specific visual encoders:
Vision Transformer (ViT) encoder, initialized from SigLIP2 [62], for image understanding. It transforms an image into approximately 4,900 tokens. Variational Autoencoder (VAE), initialized from FLUX [35], for image generation. It encodes an image into a latent grid of $\times64$, resulting in 4,096 tokens.
With these two visual encoders, BAGEL employs an expert-routing mechanism with hard gating: the understanding expert activates the ViT encoder for image understanding and text generation, while the generation expert activates the VAE decoder for high-quality image synthesis conditioned on text and images. This dual-pathway design enables fine-grained visual generation and high-level semantic grounding within a unified autoregressive modeling framework.
@@
  • Cross-Entropy Loss for text prediction:
@@
    where $x_{i}$ denotes the target text token, $\hat{x}_{i}$ is the predicted token and $C$ is the number of token classes.
  • Mean Squared Error (MSE) Loss. For denoising-based generation under the Rectified Flow [40] paradigm, given a clean latent $\mathbf{x}_{0}$ and a Gaussian noise sample $\mathbf{x}_{1}$, we construct the noisy latent $\mathbf{x}_{t}$ via linear interpolation:
@@
    The model is then trained to predict the velocity field using the MSE loss:
@@
  • The total loss is a combination of both objectives:
@@
    where ${\lambda}_{\text{CE}}$ is a coefficient to balance two losses.
@@
At inference time, BAGEL operates autoregressively over interleaved multi-modal token sequences: Understanding tasks (e.g., visual question answering, multi-modal reasoning): BAGEL consumes ViT-encoded image tokens and text tokens, and outputs the next text token via standard next-token prediction. Generation tasks (e.g., text-to-image synthesis, image editing): BAGEL predicts VAE tokens conditioned on text or image prompts via Rectified Flow process. This unified inference mechanism enables BAGEL to handle diverse multi-modal tasks within a single interface.
@@
As discussed above, while BAGEL provides a unified modeling backbone, it faces significant computational bottlenecks when applied to step-wise multi-modal reasoning. Unlike text-only CoT, where each reasoning step typically involves 300 tokens, multi-modal CoT often requires both image understanding and image generation within each step, introducing substantial overhead: generating an image via VAE incurs approximately 4,096 tokens and encoding an image via ViT for understanding introduces an additional 4,900 tokens. This results in nearly 9,000 visual tokens, in addition to around 300 text tokens, per reasoning step, leading the overall training and inference costs to be prohibitively expensive. Moreover, this token-intensive formulation imposes significant challenges on model optimization, often hindering convergence and limiting generalization, particularly in complex tasks that require long and compositional reasoning chains.
3 Method
Figure 2: Overview of the Uni-CoT framework. Uni-CoT consists of two complementary reasoning branches: (1) Macro-Level CoT, which decomposes a complex task into simpler subtasks and aggregates their outcomes to synthesize the final result. To reduce learning and computational overhead, intra-subtask reasoning is kept implicit. This process is enforced through a macro attention mask that reveals only the system prompt, high-level plans, and subtask outputs. (2) Micro-Level CoT, which executes individual subtasks while filtering out irrelevant information. It is modeled as a Markov Decision Process (MDP), where each reasoning and self-reflection step depends solely on the previous state and the current prompt. This process is enforced through a micro attention mask that restricts visibility to the last state and current instruction. High-resolution depictions of the macro and micro attention masks are shown in Figure S1.
To tackle the inherent complexity of multi-modal CoT, Uni-CoT adopts a hierarchical framework to minimize the token interactions in reasoning. As illustrated in Figure 2, the reasoning process is organized into two levels: a Macro-Level CoT, responsible for high-level task planning and summarization (Sec.3.1); and a Micro-Level CoT, which generates stable and reliable results for each subtask (Sec.3.2). This two-tiered design enables more efficient reasoning, stronger generalization, and improved interpretability across diverse multi-modal tasks. The training procedures of this framework are further detailed in Sec. 3.3.
@@ 3.1 Macro-Level CoT: Planning Strategies
Given a complex task, humans typically begin by outlining one or more abstract pathways toward the goal, rather than considering low-level details from the outset [13, 17]. Inspired by this cognitive behavior, a core function of the Macro-Level CoT is to formulate a global plan that decomposes the task into a set of simple, tractable subtasks.
@@
Specifically, drawing inspiration from [16, 77, 49, 79], we propose three global planning mechanism:
1. Sequential Decomposition: Humans often tackle complex tasks by addressing intermediate goals in a step-by-step manner. Analogously, we define a sequential decomposition strategy, wherein a task is split into a sequence of subtasks, to simplify the overall reasoning path and improve traceability. 2. Parallel Decomposition: Humans often decompose complex tasks into several independent components that can be solved concurrently, thereby accelerating the overall process. Inspired by this strategy, we propose a parallel decomposition approach that allows the model to reason over multiple subtasks simultaneously, improving both efficiency and scalability. 3. Progressive Refinement: In uncertain environments, such as navigating a maze, human often forego rigid plans and instead refine their decisions iteratively. To emulate this behavior, we propose a progressive refinement strategy, wherein the model incrementally refines its plan and revise earlier steps if inconsistencies arise, thereby supporting adaptive reasoning. These strategies empower the Macro-Level CoT to effectively address the “what-to-do” aspect of reasoning.
@@
In addition to high-level planning, the Macro-Level CoT is responsible for analyzing and synthesizing the outputs of individual subtasks into a coherent final solution, as illustrated in Figure 2. Importantly, the internal reasoning processes within each subtask are abstracted away from the macro perspective, enabling the macro planner to concentrate solely on structural decomposition and the global reasoning trajectory. To enforce this separation of concerns during both training and inference, we introduce a macro attention mask (Figure 2 or Figure S1 (a)). This mechanism selectively exposes only the system prompt, macro-level planning outputs, the outcome of each subtask, and the final result. All intermediate reasoning traces, such as textual rationales or image edits generated during micro-level execution, are fully masked. By restricting visibility to high-level signals, the Macro-Level CoT preserves a top-down reasoning perspective, thereby enhancing modularity and scalability in multi-modal CoT.
Furthermore, with these designs, training the model to learn the Macro-Level CoT is simplified into two key tasks: 1) generating global plans based on user text and image prompts, and 2) synthesizing the final results based on the outputs of the subtasks. These focused training objectives significantly stabilize and simplify the training process.